OpenStudy (anonymous):

use the method of substitution to solve the system of linear equations 7a=4-b and 7a=7-b

OpenStudy (zzr0ck3r):

if 7a = 4-b and 7a=7-b then 7a = 4-b=7-b so 4-b=7-b can you solve for b?

OpenStudy (anonymous):

no

OpenStudy (zzr0ck3r):

the system has no solutions 4-b=7-b implies 4=7

OpenStudy (anonymous):

OK thank you so much

OpenStudy (whpalmer4):

Those lines are parallel...

OpenStudy (whpalmer4):

So they never intersect, and there are no solutions.

OpenStudy (anonymous):

thank you